# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 3 Some topological constraints imposed by Maslov condition
    $Fix(G)\subset h^{-1}(0)$ up to addition of a constant to $h$, $h^{-1}(c)$ is connected for $c\in\mathbb{R}$, and
@@
    the $G$-action is semi-free.
Proof.
(i). Let $X$ be the Hamiltonian vector field of the moment map $h$ associated to the $G$-action. Observe that
$-dh=d\lambda(X_{h},\cdot)=L_{X_{h}}\lambda-d(\lambda(X_{h}))=-d(\lambda(X_{h}))$ (3)
with $L_{X_{h}}\lambda=0$ since $\lambda$ is $G$-invariant, so $\lambda(X_{h})-h$ is a constant. Up to an addition of a constant to $h$ we may assume that
$h=\lambda(X_{h})\quad\text{and hence \quad$\lambda(X_{h})=0$ on $h^{-1}(0)$},$ (4)
then
$Fix(G)\subset h^{-1}(0)$
since $\lambda(X_{h})$ is constant on level sets of $h$.
We claim that $h^{-1}(0)$ is connected. Let $\nabla h$ denote the gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metic on $M$. Then the vector fields $-\nabla h|_{h>0}$ and $\nabla h|_{h<0}$ together induces a deformation retraction from $M$ onto $h^{-1}(0)$, preserving level sets of $h$, and with $h^{-1}(0)$ fixed all the time, so $M$ is deformation equivalent and hence homotopic to $h^{-1}(0)$. Since $M$ is connected by hypothesis, so is $h^{-1}(0)$. For $c\neq 0$, $h^{-1}(c)$ is a circle bundle over the reduced space $M_{red}\cong h^{-1}(c)/G\cong h^{-1}(0)/G$. Since $h^{-1}(0)/G$ is connected, so is $h^{-1}(c)$ for $c\in\mathbb{R}$.
(ii). Suppose that $(M,\omega=d\lambda,G)$ contains exceptional $G$-orbits. Since $\omega$ is exact there are no symplectic exceptional spheres in $M$, as all critical points are on the same level set of the associated moment map $h$ say, $h^{-1}(0)$. Now assume that the $G$-action on $M$ has $n\in\mathbb{N}$ fixed points $p_{1},p_{2},\dots,p_{n}$. Maslov condition implies that in a small Darboux chart $U_{k}$ of $p_{k}$ $G$ acts as the the $S^{1}$ subgroup
$\begin{pmatrix}\cos m_{k}\theta&-\sin m_{k}\theta\\ \sin m_{k}\theta&\cos m_{k}\theta\end{pmatrix},\quad m_{k}\in\mathbb{Z}\setminus\{0\},\quad|m_{k}|>1\ \text{for some}\ 0\leq k\leq n.$
As $h^{-1}(0)$ is $G$-invariant and connected all elements in $h^{-1}(0)\setminus Fix(G)$ and hence in $M\setminus Fix(G)$ have isomorphic stabilizers, hence all $m_{k}$ are equal, the $G$-action is not effective unless either $m_{k}=1$ for all $k$, or $m_{k}=-1$ for all $k$. So there are no exceptional orbits of an effective special Hamiltonian action on $(M,\omega)$ if $\omega$ is exact. We conclude that the $G$-action on $(M,\omega=d\lambda)$ is semi-free. ∎
4 Simply-connected Exact Special Hamiltonian $S^{1}$-spaces
The rest of this paper will focus on the classification of connected exact special Hamiltonian $S^{1}$-spaces. We assume that $Fix(G)$ is a nonempty finite set as before.
Let
$\EuFrak{r}:M\to M_{red}$
denote the symplectic reduction of the $G$-action, we also denote
$\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}:h^{-1}(0)\to h^{-1}(0)/G=M_{red}$
the reduction of $\EuFrak{r}$ restricted to the level set $h^{-1}(0)$.
Proposition 4.1.
Let $(M,\omega=d\lambda,G)$ be a connected, 1-connected exact Hamiltonian $S^{1}$-space with $c_{1}(M,\omega)=0$, $\lambda$ being $G$-invariant, and $G$ acts semi-freely on $M$ with moment map $h:M\to\mathbb{R}$, and $Fix(G)\subset h^{-1}(0)$ is a finite set.
  1. (i).
    $h^{-1}(c)/G$ is diffeomorphic to a disc.
  2. (ii).
@@ 4 Simply-connected Exact Special Hamiltonian $S^{1}$-spaces
    $\lambda(X_{h})=h$ on $M$. Moreover $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. In particular, for $q\in h^{-1}(0)\setminus Fix(G)$, the symplectic normal space $T^{\omega}_{q}(h^{-1}(0)):=\{v\in T_{q}M\mid\omega_{q}(v,u)=0,\ \forall u\in T_{q}(h^{-1}(0))\}$ is generated by $(X_{h})_{q}$.
Proof.
@@
(i). Recall that $h^{-1}(c)$ is connected for all $c$. Observe that the gradient flows of ±$\pm\nabla h$ induce a deformation retract from $M$ to $h^{-1}(0)$, keeping $h^{-1}(0)$ fixed all the time. So $M$ is homotopic to $h^{-1}(0)$. Since $M$ is 1-connected, so is $h^{-1}(0)$. $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)$ is a singlar fibration over $h^{-1}(0)/G$, So $h^{-1}(0)/G$ is 1-connected. Moreover, the gradient flows of ±$\pm\nabla f$ induce a homeomorphism between the quotient spaces $h^{-1}(c)/G$ and $h^{-1}(0)/G$ for all $c\in\mathbb{R}$. So $h^{-1}(c)/G$ is 1-connected for all $c$ provided that $M$ is 1-connected.
@@
(ii). $\lambda$ is a $G$-invariant primitive 1-form of $\omega$, hence $0=L_{X_{h}}\lambda=d(\lambda(X_{h}))+d\lambda(X_{h},\cdot)=d(\lambda(X_{h}))-dh$. Since $d(\lambda(X_{h}))-dh=0$ on $h^{-1}(0)$, $\lambda(X_{h})$ is constant on $h^{-1}(0)$. $h^{-1}(0)$ is connected and $\lambda(X_{h})=0$ at $p\in Fix(G)$ so $\lambda(X_{h})=0$ and hence $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. Therefore $(X_{h})_{q}\in T^{\omega}_{q}(h^{-1}(0))$ all $q\in h^{-1}(0)\setminus Fix(G)$. Since $\dim(T^{\omega}_{q}(h^{-1}(0))+\dim T_{q}(h^{-1}(0))=\dim T_{q}M=4$, $X_{h}\subset T(h^{-1}(0))$ generates the field of symplectic normal spaces of $T(h^{-1}(0)\setminus Fix(G))$. ∎
@@
Note that Property (ii) above implies that for any embedded smooth curve $\gamma\subset M_{red}$, its preimage $L_{\gamma}:=\EuFrak{r}_{0}^{-1}(\gamma)\subset h^{-1}(0)$ is an immersed monotone Lagrangian surface of $(M,\omega)$. For example, if $\gamma$ is a circle missing the set $\EuFrak{r}_{0}(Fix(G))$ then $L_{\gamma}$ is an embedded monotone Lagrangian torus. If $\gamma$ is a circle which meets the set $\EuFrak{r}_{0}(Fix(G))$ at exactly one point, then $L_{\gamma}$ is a Lagrangian Whitney sphere, i.e., an immersed monotone Lagrangian torus with one transversal self-intersection point. If $\gamma$ is an embedded arc which connects exactly two distinct points in $\EuFrak{r}_{0}(Fix(G))$ and at its endpoint, then $L_{\gamma}$ is an embedded Lagrangian sphere.
It turns out that a 1-connected exact special Hamiltonian 4-space $(M,\omega,G)$ with finite $Fix(G)$ is symplectically the plumbing of a finite number of $T^{*}S^{2}$.
@@
Let $\gamma:(-\infty,\infty)\to h^{-1}(0)/G=M_{red}\cong\mathbb{C}$ be a smoothly embedded curve so that the complement $(h^{-1}(0)/G)\setminus\gamma$ consists of two connected components, and $\gamma(t_{i})=p_{i}$ for some $0=t_{0}<t_{1}<t_{2}<\cdots<t_{n}<\infty$, where $\{p_{0},p_{1},\dots,p_{n}\}=Fix(G)$. Let $\gamma_{i}:=\gamma([t_{i-1},t_{i}])$, $\gamma_{0}=\gamma((-\infty,0])$, $\gamma_{n+1}=\gamma([t_{n},\infty))$. Let $S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i})\subset h^{-1}(0)$, $S_{i}$ is a $G$-invariant Lagrangian sphere which is smooth except perhaps at the two points $p_{i-1}$ and $p_{i}$.
Let $U_{i}\subset M$ be a $G$-invariant symplectic Darboux chart centered at $p_{i}$, $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$. Let $U:=\cup_{i=0}^{n}U_{i}$. By shrinking $U_{i}$ if necessary we may assume that $=\textsf{G}$ G on $U$. Then $=\textsf{h}$ h and $X_{h}=X_{\textsf{h}}$ on $U$. Each $U_{i}$ is symplectically identified with a open neighborhood of $0\in\mathbb{C}^{2}$ with $p_{i}$ identified with the origin $0\in\mathbb{C}^{2}$. Recall the symplectic reduction $\EuFrak{r}:M\to M_{red}\cong\mathbb{C}$ of the $G$-action and its restriction $\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}$ to $h^{-1}(0)$. We may assume that $\EuFrak{r}(z_{1},z_{2})=\frac{1}{2}(z_{1}^{2}+z_{2}^{2})+\gamma(t_{i})$ when restricted to each $U_{i}$, and, by perturbing $\gamma$ if necessary, that $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i-1}\cap h^{-1}(0))=\{y_{1}=0=y_{2}\}\cap U_{i-1}$ and $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i}\cap h^{-1}(0))=\{x_{1}=0=x_{2}\}\cap U_{i}$ are Lagrangian discs for each $i$. Then
@@
$S_{i}$ can be oriented so that the intersection number is $S_{i}\cdot S_{i+1}=1$ for $i=1,...,n-1$. The union $\mathcal{S}:=\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of Lagrangian spheres.
By Lagrangian Neighborhood Theorem [13, 15], a tubular neighborhood $V_{i}\subset M$ of the Lagrangian sphere $S_{i}$ is symplectomorphic to a tubular neighborhood of the 0-section of the cotangent bundle $(T^{*}S_{i},\omega_{can})$ of $S_{i}$ equipped with the canonical symplectic structure. $T^{*}S_{i}$ and $T^{*}S_{i+1}$ are plumbed together near $p_{i}$ so that tangent and cotangent spaces of $S_{i}$ are identified with the cotangent and tangent spaces of $S_{i+1}$ respectively for $i=1,...,n-1$:
@@
[…]
Indeed we can take an open cover of $\EuFrak{r}(M)=h^{-1}(0)/G\cong\mathbb{R}^{2}$ by 1-connected domains $V_{i}$, $i=1,\cdots,n$ so that $\EuFrak{r}(M)=\cup_{i}V_{i}$, $S_{i}\subset V_{i}$, $V_{i}\cap V_{j}=\emptyset$ if $|i-j|>1$, and $V_{i}\cap V_{i+1}$ is 1-connected. Then each $\EuFrak{r}^{-1}(V_{i})$ is symplectomorphic to an open neighborhood in $T^{*}S_{i}$ of the zero section $S_{i}$ of $T^{*}S_{i}$, and $V_{i}\cap V_{i+1}$ is where the plumbing of $T^{*}S_{i}$ and $T^{*}S_{i+1}$ takes place. This completes the proof of Theorem 1.2.
@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
The G-action on $\mathbb{C}^{2}$ can be lifted to a special Hamiltonian $S^{1}$-action on $\mathbb{C}^{3}$ which acts trivially on the $z_{3}$-coordinate:
$\tilde{\textsf{G}}:=\Big{\{}\tilde{\textsf{G}}(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\mid\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}\subset SU(3).$
The action of $\tilde{\textsf{G}}$ preserves $\mathcal{W}_{n}$ and restricts to a special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$, denoted as G for simplicity, with moment map
@@
$Fix(\textsf{G})=\{p_{k}:=(0,0,\xi^{k})|k=0,...,n\}\subset\textsf{h}^{-1}(0),\quad\xi:=e^{\frac{2\pi i}{n+1}}.$
The tangent space of $\mathcal{W}_{n}$ at $p_{k}$ is $T_{p_{k}}\mathcal{W}_{n}=\mathbb{C}^{2}_{z_{1}z_{2}}$ for each $k$. Since h is independent of $z_{3}$, the Morse index of h at $p_{k}$ is $2$ as in Example 2.2. It follows that the G-action on $\mathcal{W}_{n}$ is special Hamiltonian. This completes the proof of Proposition 1.3.
@@
[…]
Consider the level set $\textsf{h}^{-1}(0)$. Without loss of generality, we may restrict to a G-invariant Stein domain
@@
for some $c>0$ and $D=\{|z|\leq r\}$ for some $r>1$.
@@
which contains $Fix(\textsf{G})$ and is transversal to G-orbits except at $Fix(\textsf{G})$. Let $\Sigma_{\theta}:=\textsf{G}(\theta)(\Sigma)$, then $\Sigma_{\theta+\pi}=\Sigma_{\theta}$ for each $\theta$, and $\Sigma_{\theta}\cap\Sigma_{\theta^{\prime}}=Fix(\textsf{G})$ if $\theta\neq\theta^{\prime}\mod\pi$. Moreover,
@@
For each $\theta$ the symplectic reduction map $\EuFrak{r}_{0}|_{\Sigma_{\theta}}:\Sigma_{\theta}\to\mathbb{C}$ is a $2:1$ branched covering map branching at $Fix(\textsf{G})$, with branching locus $\{\xi\in\mathbb{C}\mid\xi^{n+1}=1\}\subset\mathbb{C}$. Let $g$ and $b$ be the genus and boundary components of $\Sigma$ (hence each $\Sigma_{\theta}$), then $2-2g-b=2-(n+1)$. Observe that the subgroup $K:=\{\pm Id\}\subset\textsf{G}$ acts on $\partial\Sigma$, $\partial\Sigma/K$ is a simple closed circle which is diffeomorphic to $\partial D$ under the map $\EuFrak{r}_{0}$. Therefore $b=1$ or $2$, $\Sigma$ is a surface of genus $g=\lfloor\frac{n}{2}\rfloor$ with $b=(n+1)-2g$ boundary components. More explicitly,
$(g,b)=\begin{cases}(m,1)&\text{ if $n=2m\geq 0$ is even},\\ (m,2)&\text{ if $n=2m+1\geq 1$ is odd}.\end{cases}$
@@
$\gamma_{k}=\Big{\{}e^{i\theta}\in\mathbb{C}_{z_{3}}\Big{|}\frac{2(k-1)\pi}{n+1}\leq\theta\leq\frac{2k\pi}{n+1}\Big{\}},\quad k=1,\dots,n,$ (5)
@@
$\gamma_{k}$ is an arc between $\xi^{k-1}$ and $\xi^{k}$. Orient $\gamma_{k}$ so that its tangents point from $\xi^{k-1}$ to $\xi^{k}$. Then $\textsf{S}_{k}:=(\EuFrak{r}_{0})^{-1}(\gamma_{k})$ is a G-invariant Lagrangian sphere, $\hat{\gamma}_{k}:=\textsf{S}_{k}\cap\Sigma_{0}$ a simple closed curve, and $\textsf{S}_{k}\cap\Sigma_{\theta}=\textsf{G}(\theta)(\hat{\gamma}_{k})$, $\textsf{S}_{k}=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\textsf{G}(\theta)(\hat{\gamma}_{k})$. The union $\cup_{k=1}^{n}\textsf{S}_{k}$ is an $A_{n}$-string of Lagrangian spheres. $\mathcal{W}_{n}$ is an $A_{n}$-surface singularity as in [16].
Proof of Theorem 1.4.

### v4

@@ 4 Simply-connected Exact Simple SHam1-spaces
    $h^{-1}(c)/\Phi$ is 1-connected hence homeomorphic to a disc.
@@
    $\lambda(X_{h})=h$ on $M$. Moreover $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. In particular, for $q\in h^{-1}(0)\setminus Fix(\Phi)$, the symplectic normal space $T^{\omega}_{q}(h^{-1}(0)):=\{v\in T_{q}M\mid\omega_{q}(v,u)=0,\ \forall u\in T_{q}(h^{-1}(0))\}$ is generated by $(X_{h})_{q}$.
@@
Proof. (i) Recall that $h^{-1}(c)$ is connected for all $c$. Observe that the gradient flows of $\pm\nabla h$ induce a deformation retract from $M$ to $h^{-1}(0)$, keeping $h^{-1}(0)$ fixed all the time. So $M$ is homotopic to $h^{-1}(0)$. Since $M$ is 1-connected, so is $h^{-1}(0)$. $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/\Phi$ is a singlar fibration over $h^{-1}(0)/\Phi$, hence $h^{-1}(0)/\Phi$ is 1-connected. Moreover, the gradient flows of $\pm\nabla h$ induce a homeomorphism between the quotient spaces $h^{-1}(c)/\Phi$ and $h^{-1}(0)/\Phi$ for all $c\in\mathbb{R}$, so $h^{-1}(c)/\Phi$ is 1-connected for all $c$ provided that $M$ is 1-connected.
@@
(ii) $\lambda$ is a $\Phi$-invariant primitive 1-form of $\omega$, hence $0=L_{X_{h}}\lambda=d(\lambda(X_{h}))+d\lambda(X_{h},\cdot)=d(\lambda(X_{h}))-dh$. Since $d(\lambda(X_{h}))-dh=0$ on $h^{-1}(0)$, $\lambda(X_{h})$ is constant on $h^{-1}(0)$. $h^{-1}(0)$ is connected and $\lambda(X_{h})=0$ at $p\in Fix(\Phi)$ so $\lambda(X_{h})=0$ and hence $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. Therefore $(X_{h})_{q}\in T^{\omega}_{q}(h^{-1}(0))$ all $q\in h^{-1}(0)\setminus Fix(\Phi)$. Since $\dim(T^{\omega}_{q}(h^{-1}(0))+\dim T_{q}(h^{-1}(0))=\dim T_{q}M=4$, $X_{h}\subset T(h^{-1}(0))$ generates the field of symplectic normal spaces of $T(h^{-1}(0)\setminus Fix(\Phi))$. $\Box$
@@
Property (ii) above implies that for any embedded smooth curve $\gamma\subset M_{0}/\Phi$, its preimage $L_{\gamma}:=\EuFrak{r}_{0}^{-1}(\gamma)\subset h^{-1}(0)$ is an immersed monotone Lagrangian surface of $(M,\omega)$. For example, if $\gamma$ is a circle missing the set $(Fix(\Phi)$ then $L_{\gamma}$ is an embedded monotone Lagrangian torus. If $\gamma$ is a circle which meets the set $\EuFrak{r}_{0}(Fix(\Phi))$ at exactly one point, then $L_{\gamma}$ is a Lagrangian Whitney sphere, i.e., an immersed monotone Lagrangian torus with one transversal self-intersection point. If $\gamma$ is an embedded arc which connects exactly two distinct points in $\EuFrak{r}_{0}(Fix(\Phi))$ and at its endpoint, then $L_{\gamma}$ is an embedded Lagrangian sphere.
It turns out that a 1-connected exact SHam1-space $(M,\omega,\Phi)$ with finite $Fix(\Phi)$ is symplectically the plumbing of a finite number of $T^{*}S^{2}$.
@@
Let $\gamma:(-\infty,\infty)\to h^{-1}(0)/\Phi\cong\mathbb{C}$ be a smoothly embedded curve so that the complement $(h^{-1}(0)/\Phi)\setminus\gamma$ consists of two connected components, and $\gamma(t_{i})=p_{i}$ for some $0=t_{0}<t_{1}<t_{2}<\cdots<t_{n}<\infty$, where $\{p_{0},p_{1},\dots,p_{n}\}=Fix(\Phi)$. Let $\gamma_{i}:=\gamma([t_{i-1},t_{i}])$, $\gamma_{0}=\gamma((-\infty,0])$, $\gamma_{n+1}=\gamma([t_{n},\infty))$. Let $S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i})\subset h^{-1}(0)$, then by Proposition 4.1 (ii) $S_{i}$ is a $\Phi$-invariant Lagrangian sphere which is smooth except perhaps at the two points $p_{i-1}$ and $p_{i}$.
Let $U_{i}\subset M$ be a $\Phi$-invariant symplectic Darboux chart centered at $p_{i}$, $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$. Let $U:=\cup_{i=0}^{n}U_{i}$. By shrinking $U_{i}$ if necessary we may assume that $\Phi=\Theta$ on $U$. Then $h=\textsf{h}$ and $X_{h}=X_{\textsf{h}}$ on $U$. Each $U_{i}$ is symplectically identified with a open neighborhood of $0\in\mathbb{C}^{2}$ with $p_{i}$ identified with the origin $0\in\mathbb{C}^{2}$. Recall $\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}/\Phi$. We may assume that $\EuFrak{r}(z_{1},z_{2})=\frac{1}{2}(z_{1}^{2}+z_{2}^{2})+\gamma(t_{i})$ when restricted to each $U_{i}$, and, by perturbing $\gamma$ if necessary, that $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i-1}\cap h^{-1}(0))=\{y_{1}=0=y_{2}\}\cap U_{i-1}$ and $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i}\cap h^{-1}(0))=\{x_{1}=0=x_{2}\}\cap U_{i}$ are Lagrangian discs for each $i$. Then
@@
$S_{i}$ can be oriented so that the intersection number is $S_{i}\cdot S_{i+1}=1$ for i=1,,n1i=1,...,n-1. The union $\mathcal{S}:=\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of Lagrangian spheres.
By Lagrangian Neighborhood Theorem [13, 15], a tubular neighborhood $V_{i}\subset M$ of the Lagrangian sphere $S_{i}$ is symplectomorphic to a tubular neighborhood of the 0-section of the cotangent bundle $(T^{*}S_{i},\omega_{can})$ of $S_{i}$ equipped with the canonical symplectic structure. $T^{*}S_{i}$ and $T^{*}S_{i+1}$ are plumbed together near $p_{i}$ so that tangent and cotangent spaces of $S_{i}$ are identified with the cotangent and tangent spaces of $S_{i+1}$ respectively for i=1,,n1i=1,...,n-1:
@@
where $x=(x_{1},x_{2})$ are coordinates of $S_{i}$ near $p_{i}$, $y=(y_{1},y_{2})$ are the fiber coordinates of the cotangent space $T^{*}_{p_{i}}S_{i}$.
Indeed we can take an open cover of $\EuFrak{r}_{0}=h^{-1}(0)/\Phi\cong\mathbb{R}^{2}$ by 1-connected domains $V_{i}\subset\EuFrak{r}_{0}$, $i=1,\cdots,n$ so that $\EuFrak{r}_{0}=\cup_{i}V_{i}$, $\gamma_{i}\subset V_{i}$, $V_{i}\cap V_{j}=\emptyset$ if |ij|>1|i-j|>1, and $V_{i}\cap V_{i+1}$ is 1-connected. Then each $\EuFrak{r}^{-1}(V_{i})$ is symplectomorphic to an open neighborhood in $T^{*}S_{i}$ of the zero section $S_{i}$ of $T^{*}S_{i}$, and $V_{i}\cap V_{i+1}$ is where the plumbing of $T^{*}S_{i}$ and $T^{*}S_{i+1}$ takes place. This completes the proof of Theorem 1.3 if $|Fix(\Phi)|=n+1\geq 2$.
If $|Fix(\Phi)|=1$ then $M$ is diffeomorphic to a symplectic 4-space, and we may take the fixed point to be $0\in\mathbb{R}^{4}$. $\Box$
@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
The standard SHam1-action $\Theta$ on $\mathbb{C}^{2}$ can be lifted to a SHam1-action on $\mathbb{C}^{3}$ which acts trivially on the $z_{3}$-coordinate:
$\tilde{\Theta}:=\Big\{\tilde{\Theta}_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\mid\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big\}\subset SU(3).$
The action of $\tilde{\Theta}$ preserves $\mathcal{W}_{n}$ and restricts to a special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$, denoted as $\Theta$ for simplicity, with moment map
@@
$Fix(\Theta)=\{p_{k}:=(0,0,\xi^{k})|k=0,...,n\}\subset\textsf{h}^{-1}(0),\quad\xi:=e^{\frac{2\pi i}{n+1}}.$
The tangent space of $\mathcal{W}_{n}$ at $p_{k}$ is $T_{p_{k}}\mathcal{W}_{n}=\mathbb{C}^{2}_{z_{1}z_{2}}$ for each $k$. Since h is independent of $z_{3}$, the Morse index of h at $p_{k}$ is $2$ as in Example 2.2. It follows that the $\Theta$-action on $\mathcal{W}_{n}$ is special Hamiltonian. This completes the proof of Proposition 1.4.
Remark 5.1.
(Equivarience between the $\Theta$-action and its inverse $\Theta^{-1}$ on $(\mathcal{W}_{n},\omega_{n})$.)
On $(\mathcal{W}_{n},\omega_{n})$ consider the symplectic diffeomorphism $f:\mathcal{W}_{n}\to\mathcal{W}_{n}\subset\mathbb{C}^{3}$
$f(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})=(-x_{1},-y_{1},x_{2},y_{2},x_{3},y_{3}).$
$f$ can be represented by the unitary matrix
$\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\quad f^{-1}=f.$
Then
$f^{-1}\circ\Theta^{-1}\circ f=\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$\begin{bmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{bmatrix}=\Theta.$
So $\Theta$ and $\Theta^{-1}$ are symplectically equivariant SHam-1 actions on $\mathcal{W}_{n}$. $\Box$
@@
Observe that the value of $z_{3}$ is constant along any integral curve $\ell$ of $\nabla\textsf{h}$ and hence on the cylinder $Orb_{\Theta}(\ell)$ degenerated or not. So for $z\in\mathbb{C}_{z_{3}}$ the preimage $C_{z}:=(\pi^{\prime})^{-1}(z)$ is a $\Theta$-invariant cylinder degenerate only at point $p_{k}=(0,0,\xi^{k})$ if $z=\xi^{k}$ for some $k=0,1,\dots,n$, with tangent space $T_{p}C_{z}=span\{X,\nabla\textsf{h}\}$ when $p\neq p_{k}$.
Consider the level set $\textsf{h}^{-1}(0)$. Without loss of generality, we may restrict to a $\Theta$-invariant Stein domain
@@
for some $c>0$, $D=\{|z|\leq r\}$ for some $r>1$ and $Fix(\Theta)\subset D$.
@@
which contains $Fix(\Theta)$ and is transversal to $\Theta$-orbits except at $Fix(\Theta)$. Let $\Sigma_{\theta}:=\Theta_{\theta}(\Sigma)$, then $\Sigma_{\theta+\pi}=\Sigma_{\theta}$ for each $\theta$, and $\Sigma_{\theta}\cap\Sigma_{\theta^{\prime}}=Fix(\Theta)$ if $\theta\neq\theta^{\prime}\mod\pi$. Moreover,
@@
For each $\theta$ the symplectic reduction map $\EuFrak{r}_{0}|_{\Sigma_{\theta}}:\Sigma_{\theta}\to\mathbb{C}$ is a $2:1$ branched covering map branching at $Fix(\Theta)$, with branching locus $\{\xi\in\mathbb{C}\mid\xi^{n+1}=1\}\subset\mathbb{C}$. Let $g$ and $b$ be the genus and boundary components of $\Sigma$ (hence each $\Sigma_{\theta}$), then 22gb=2(n+1)2-2g-b=2-(n+1). Observe that the subgroup $K:=\{\pm Id\}\subset\Theta$ acts on $\partial\Sigma$, $\partial\Sigma/K$ is a simple closed circle which is diffeomorphic to $\partial D$ under the map $\EuFrak{r}_{0}$. Therefore $b=1$ or $2$, $\Sigma$ is a surface of genus $g=\lfloor\frac{n}{2}\rfloor$ with b=(n+1)2gb=(n+1)-2g boundary components. More explicitly,
[…]
@@
(4) $\gamma_{k}=\Big\{e^{i\theta}\in\mathbb{C}_{z_{3}}\Big|\frac{2(k-1)\pi}{n+1}\leq\theta\leq\frac{2k\pi}{n+1}\Big\},\quad k=1,\dots,n,$
@@
$\gamma_{k}$ is an arc between $\xi^{k-1}$ and $\xi^{k}$. Orient $\gamma_{k}$ so that its tangents point from $\xi^{k-1}$ to $\xi^{k}$. Then $\textsf{S}_{k}:=(\EuFrak{r}_{0})^{-1}(\gamma_{k})$ is a $\Theta$-invariant Lagrangian sphere, $\hat{\gamma}_{k}:=\textsf{S}_{k}\cap\Sigma_{0}$ a simple closed curve, and $\textsf{S}_{k}\cap\Sigma_{\theta}=\Theta_{\theta}(\hat{\gamma}_{k})$, $\textsf{S}_{k}=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\Theta_{\theta}(\hat{\gamma}_{k})$. The union $\cup_{k=1}^{n}\textsf{S}_{k}$ is an $A_{n}$-string of Lagrangian spheres. $\mathcal{W}_{n}$ is an $A_{n}$-surface singularity as in [16].
Proof of Theorem 1.8